A small jet airplane has a total wing area of 77.5 m2 and a mass of 7.03 x 104 kg (a) If this jet is in horizontal flight, determine the pressure difference between the lower and upper surfaces of the wings Pa (b) When the speed of air traveling over the wing is 237 m/s, determine the speed of air under the wing. Use 1.29 kg/m3 as the density of air. m/s (c) Why do all aircraft have a maximum operational altitude? O The density of air decreases with higher altitude, which increases O The density of air increases with higher altitude, which decreases O The density of air increases with higher altitude, which increases the O The density of air decreases with higher altitude, which decreases the pressure difference until it cannot support the aircraft. the pressure difference until it cannot support the aircraft. pressure difference until it cannot support the aircraft. the pressure difference until it cannot support the aircraft.
Answer #1

Solution)

Part a)

Weight of the plane ,W = m * g

= 7.03*10^4*9.8

= 688940 N

Now, pressure P = F / A

And force equal to the weight of the plane

So, P = W / A

= 688940 / 77.5

= 8889.54 Pa (Ans)

====

b)

From Bernoulli's equation:

p1 + 1/2*rho*v1^2 = p2 + 1/2*rho*v2^2

for points 1 and 2 along the same path in steady, incompressible flow.

pt + 1/2*rho*vt^2 = pb + 1/2*rho*vb^2

where t = top

And b = bottom

So,

1/2*rho*vt^2 - 1/2*rho*vb^2 = pb - pt

0.5 * 1.29 * (237)^2 - 0.5 * 1.29 * vb^2 = 8889.54

36229.005- 0.645 vb^2 = 8889.54

27339.46= 0.645 vb^2

vb=205.88 m/s (Ans)

======

Good luck!:)